A study was designed to compare the attitudes of two groups of nursing students towards computers. Group 1 had previously taken a statistical methods course that involved significant computer interaction. Group 2 had taken a statistic methods course that did not use computers. The students' attitudes were measured by administering the Computer Anxiety Rating Scale (CARS). A random sample of 18 nursing students from Group 1 resulted in a mean score of 44 with a standard deviation of 6.5. A random sample of 11 nursing students from Group 2 resulted in a mean score of 50.2 with a standard deviation of 4.9. Can you conclude that the mean score for Group 1 is significantly lower than the mean score for Group 2? Let μ1 represent the mean score for Group 1 and μ2 represent the mean score for Group 2. Use a significance level of α=0.05 for the test. Assume that the population variances are equal and that the two populations are normally distributed.

Step 1 of 4: State the null and alternative hypotheses for the test.

Step 2 of 4: Compute the value of the t test statistic. Round your answer to three decimal places

Step 3 of 4: Determine the decision rule for rejecting the null hypothesis H0. Round your answer to three decimal places.

Step 4 of 4: State the test's conclusion.

.
Answer #1

1)

Below are the null and alternative Hypothesis,
Null Hypothesis, H0: μ1 = μ2
Alternative Hypothesis, Ha: μ1 < μ2

2)

Pooled Variance
sp = sqrt((((n1 - 1)*s1^2 + (n2 - 1)*s2^2)/(n1 + n2 - 2))*(1/n1 + 1/n2))
sp = sqrt((((18 - 1)*6.5^2 + (11 - 1)*4.9^2)/(18 + 11 - 2))*(1/18 + 1/11))
sp = 2.2801

Test statistic,
t = (x1bar - x2bar)/sp
t = (44 - 50.2)/2.2801
t = -2.719

3)


P-value Approach
P-value = 0.006
As P-value <= 0.05, reject null hypothesis.

4)

There is sufficient evidence to conclude that the mean score for Group 1 is significantly lower than the mean score for Group 2